the lone star shipping company charges$14 plus $2 a pound to ship an overnight package. discount shipping company charges $20 plus $1.50 a pound over night package. for what weight is the charge the same for the companies
Answer by ReadingBoosters(3246)

Just in case you are yet in need of a solution...
Lone Star
c = 14 + 2x, where x is the number of pounds
Discount
c = 20 + 1.5x
...
At what weight are the charges the same? Set them equal to each other and solve for x.
14 + 2x = 20 + 1.5x
Group like terms
2x - 1.5x = 20 - 14
Combine like terms
.5x = 6
Divide both sides by .5
x = 12
...
When the weight is 12 lbs, the charges at both companies are the same.
........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com